# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v1

@@ 1.1. Problem formulation and main results
Let $(\mathcal{X},d)$ be a complete separable metric space and $\mathcal{B}(\mathcal{X})$ be the $\sigma$-algebra of Borel sets of $\mathcal{X}$. Let $P(\mathcal{X})$ denote the set of Borel probability measures on $\mathcal{X}$ and $P_{p}(\mathcal{X})\subset P(\mathcal{X})$ denote the subset of measures with finite $p$-th moment for $p\geq 1$. Consider a prescribed family of probability measures $(\mu_{t})\coloneqq(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ indexed by $t$ in a time interval $I\coloneqq[0,T]\subset\mathbb{R}$. In many applications, it is of interest to have a stochastic process $(X_{t})$ on a suitable path space $\Gamma_{T}\subset\mathcal{X}^{[0,T]}$ whose one-dimensional time marginals coincide with $(\mu_{t})$. We know that this requirement does not uniquely determine the process in general. In this paper, we are interested in processes whose path laws, denoted by $\pi\in P(\Gamma_{T})$ and referred to as lifts, carry further intrinsic information about $(\mu_{t})$, as a result of a minimization problem analogous to Kantorovich’s optimal transport problem (1942) [11]. Classical choices for $\Gamma_{T}$ are the space of continuous paths and the space of càdlàg paths. Throughout this paper, we consider $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ endowed with the $\sigma$-algebra $\mathcal{C}$ generated by the corresponding evaluation maps $\{e_{t}:t\in I\}$. We also equip $\Gamma_{T}$ with the supremum distance and recall that the Borel $\sigma$-algebra generated by the corresponding open sets coincides with $\mathcal{C}$. We formulate the problem mentioned above as a variational problem on the space of path measures on $\Gamma_{T}$:
@@
$\mathrm{Lift}(\mu_{t})\coloneqq\Big{\{}\pi\in P(\Gamma_{T}):\quad(e_{t})_{\#}\pi=\mu_{t}\text{ for all }t\in I\Big{\}}.$ (1.2)
Here, $e_{t}:\Gamma_{T}\to\mathcal{X}$ is the evaluation map defined by $e_{t}(\gamma)\coloneqq\gamma_{t}$ for $\gamma\in\Gamma_{T}$.
This extends Kantorovich’s classical problem, with two fixed marginals $\mu,\nu\in P(\mathcal{X})$, to a time-dependent setting with infinitely many fixed marginals $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$. In contrast to the set of couplings $\mathrm{Cpl}(\mu,\nu)\subset P(\mathcal{X}^{2})$, which is always non-empty, the set of liftings $\mathrm{Lift}(\mu_{t})\subset P(\mathcal{X}^{I})$ can be empty. In this case the infimum (1.1) is +$\infty$+ ∞ by the usual convention.
@@
The reason for adopting a larger path space in the last case, namely the space of càdlàg curves $D(I;\mathcal{X})$, is that even $1$-absolutely continuous curves in 1-Wasserstein spaces, unlike the case $p>1$, cannot generally be lifted to measures on continuous paths. This can also be noticed through the Kolmogorov–Chentsov continuity criterion, where $p=1$ is excluded. In this work, we consider $p>1$ and study continuous Wasserstein curves with lower regularity than above.
@@
As the next step, we start from a path measure $\pi$ of finite $W^{\alpha,p}$-energy and show that its curve of one-dimensional time marginals $t\mapsto\mu_{t}\coloneqq(e_{t})_{\#}\pi$ inherits the same kind of regularity, and, moreover, its regularity is bounded from above by the energy of $\pi$. Below is Theorem 3.5.
@@
In fact, this transfer of regularity from the path measure to its curve of marginals holds for all lower-semi continuous norms listed in Table 1, as shown in Section 3.2, but is not reported here. By combining Proposition 1.1, Theorem 1.2, and the observation on $|\cdot|_{W^{\alpha,p}}$ in Table 1, we immediately draw the following conclusion as the first main result:
@@
    Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ can form a compatible collection under suitable conditions on their covariance matrices. See [18, page 49].
@@
    For more nontrivial compatible examples, see [7, Proposition 4.1] and [18, Section 2.3.2].
@@
    Construct the path measure $\pi_{n}\coloneqq(\ell)_{\#}\Upsilon_{n}\in P(\Gamma_{1})$, where $\ell:\mathcal{X}^{2^{n}+1}\to\Gamma_{1}$ is the geodesic selection and interpolation map connecting the points with constant-speed geodesics.
@@
Notice that $\left(1-2^{-(p-\alpha p)}\right)\in(0,1)$ on the parameter range above. This characterization allows us to give the result below, which is a generalization of the metric Benamou–Brenier formula [5], [4, Eq. (8.0.3)], [13, Corollary 4], to the fractional setting. Below is Corollary 3.13, where $Geo([0,1];\mathcal{X})$ denotes the set of constant-speed geodesics.
@@ 1.2. Motivating example
In the first case of this example, we revisit the well-known relationship between the continuity equation and absolutely continuous Wasserstein curves under the $p$-integrability condition, [4, Chapter 8] and [13, Section 1]. In the second case, we assume a weaker integrability condition and highlight the estimates applied in the fractional setting.
@@
This can be used to infer that $(\mu_{t})$ is an absolutely continuous curve. Moreover, as shown in [13, Theorem 4] by the Lebesgue differentiation theorem, we obtain an upper bound for the metric speed $|\dot{\mu}_{t}|\leq\|\dot{\gamma}_{t}\|_{L^{p}(\pi)}$ for a.e. $t\in I$. Raising it to power $p$ and integrating in time yields
@@ 1.4. Acknowledgements
The author would like to thank his supervisor, Matthias Erbar, for invaluable guidance and numerous discussions. He also thanks Timo Schultz for insightful conversations, which significantly improved the results and eventually led to Examples 4.3, 4.4 and 4.5. He also thanks Zhenhao Li for helpful comments on some technical details. Additionally, he extends his thanks to Kohei Suzuki and Vitalii Konarovskyi for insightful discussions.
@@ 2.1. Continuous paths on metric spaces
The Arzelà-Ascoli theorem provides a characterization of relative compactness in $\Gamma_{T}$ (see, e.g., [6, Theorem 7.2] or [9, Theorem 1.4]).
@@ 2.2. Hölder paths on metric spaces
Below, a simple yet useful characterization of Hölder continuity is provided. It states that for a continuous curve $X:[0,1]\to\mathcal{X}$ to be Hölder, it is enough to verify the Hölder condition only at the dyadic time points of $[0,1]$. The result is taken from [17, Lemma 2], whose proof is inspired by the classical proof of Kolmogorov–Čentsov continuity theorem (see e.g. [12, Theorem 2.8]). The Kolmogorov–Chentsov continuity theorem can also be proved using fractional Sobolev norms (see e.g. the proof of [21] and [9, Theorem A.10]).
Theorem 2.7 (A discrete characterization of $C^{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$ [17, Lemma 2]).
@@
While one direction is trivial, we stated the theorem in this way for its resemblance to the forthcoming Theorem 2.17. It is important to note that in the second condition, we evaluate $X$ only at a countable dense subset of the interval $[0,1]$. Thus, continuity needs to be added as an extra assumption. The result above differs slightly from the original statement in [17], where a (unique) continuous modification is constructed. Here due to the continuity assumption, the continuous modification coincides with the original curve at all times.
We recall some properties of the Hölder regularity. For below, see e.g. [9, Lemma 5.12].
@@ 2.3. $p$-variation paths on metric spaces
Two subsequent lemmas are taken from [9, Proposition 5.6].
@@
To distinguish these norms, we refer to the first one as variation, while the second one as infinitesimal variation, emphasizing that they are not necessarily identical. This is only the case for $p=1$ (see e.g. [9, Proposition 1.14]):
@@
As a next result, we recall the lower semi-continuity of $p$-variation, as stated in [9, Lemma 5.12].
@@
Let $(\mathcal{X},d)$ be a geodesic space. Take $p>1$ and let $\gamma$ be a continuous path such that $0<|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}<\infty$. Let $(D_{n})_{n\in\mathbb{N}}$ be a sequence of dissections of $[0,T]$ with shrinking mesh size $|D_{n}|\to 0$. Define $\gamma^{D_{n}}$ to be the piecewise geodesic approximation of $\gamma$ (i.e. the curve that coincides with $\gamma$ at points of $D_{n}$ and connects in between with geodesics). We know that $\gamma^{D_{n}}\to\gamma$ uniformly (see e.g. [9, Lemma 5.19]). For each $n$, the path $\gamma^{D_{n}}$ is of bounded $1$-variation, which implies, by Lemma 2.12, that $|\gamma^{D_{n}}|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}=0$ for all $p>1$. Consequently,
@@ 2.4. Fractional Sobolev paths on metric spaces
The fractional Sobolev spaces $W^{\alpha,p}$ (also known as Sobolev–Slobodeckij spaces) can be viewed as a specific instance of the broader class of Besov spaces $B^{\alpha,p,q}$ (see e.g. [15, Equation (2.1)]). In particular, when $q=p$, we have $B^{\alpha,p,p}=W^{\alpha,p}$. While the classical definition of $B^{\alpha,p,q}$ typically involves double integrals similar to the one mentioned above, there are also other function spaces whose definition involves double sums. Recently, [15] has proved that these definitions are in fact equivalent. For the purpose of our work, we only give their result for the case $q=p$:
Theorem 2.17 (A discrete characterization of $W^{\alpha,p}$ [15, Theorem 2.2]).
@@
This result is reminiscent of the discrete characterization of Hölder curves provided by [17, Lemma 2], as stated in Theorem 2.7. In the second definition, we are once again evaluating $X$ only at dyadic points (i.e. a countable subset of $[0,1]$) but this time, in the form of a double sum. Thus, continuity must be an additional assumption. This is, however, not needed for the first definition. In fact, under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the finiteness of $|X|_{W^{\alpha,p}}$ automatically implies the continuity of $X$ (see Theorem 2.19). We repeatedly rely on the equivalence result above. We proceed with the following simple observation:
@@
The estimate (2.29) below can be derived by applying the Garsia–Rodemich–Rumsey inequality, which was originally introduced in [10] for the case $\mathcal{X}=\mathbb{R}$. The proof in the general metric setting can be found in [9, Theorem A.1]). These results are well-known [9, Corollary A.2-3]
@@
Note that from (2.29), we can immediately conclude that $X$ has finite $({\alpha-1/p})^{-1}$-variation. But, as shown in [8, Theorem 2], by applying Hölder’s inequality with a clever choice of exponents, an even stronger statement can be made, namely, $X$ has finite $({\alpha})^{-1}$-variation. A consequence of this embedding is that under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the elements in $W^{\alpha,p}$ have zero infinitesimal $p$-variation by Lemma 2.12:
@@ 2.5. Sobolev paths on metric spaces
On $\mathbb{R}^{\mathrm{\,d}}$, it is well-known that the function spaces mentioned above coincide when $p>1$ (see e.g. [9, Proposition 1.45]). This result also holds in the general metric setting. For a proof in the case $p=2$, see [2, Theorem 10.2], which can be generalized to any $p>1$ and is stated below:
@@
which implies the embeddings [9, Theorem 1.47]
@@ 2.6. Geodesics on metric spaces
$d(\gamma_{s},\gamma_{t})=|t-s|d(\gamma_{0},\gamma_{1}),\quad\forall t,s\in[0,1].$

### v3

@@ 1.1. Problem formulation and main results
Let $(\mathcal{X},d)$ be a complete separable metric space and $\mathcal{B}(\mathcal{X})$ be the $\sigma$-algebra of Borel sets of $\mathcal{X}$. Let $P(\mathcal{X})$ denote the set of Borel probability measures on $\mathcal{X}$ and $P_{p}(\mathcal{X})\subset P(\mathcal{X})$ denote the subset of measures with finite $p$-th moment for $p\geq 1$. Consider a prescribed family of probability measures $(\mu_{t})\coloneqq(\mu_{t})_{t\in I}\subset P(\mathcal{X})$ indexed by $t$ in a time interval $I\coloneqq[0,T]\subset\mathbb{R}$. In many applications, it is of interest to have a stochastic process $(X_{t})$ on a suitable path space $\Gamma_{T}\subset\mathcal{X}^{[0,T]}$ whose one-dimensional time marginals coincide with $(\mu_{t})$. We know that this requirement does not uniquely determine the process in general. In this paper, we are interested in processes whose path laws, denoted by $\pi\in P(\Gamma_{T})$ and referred to as lifts, carry further intrinsic information about $(\mu_{t})$, as a result of a minimization problem analogous to Kantorovich’s optimal transport problem (1942) [12]. Typical choices for $\Gamma_{T}$ are the space of continuous paths and the space of càdlàg paths. Throughout this paper, we consider $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ endowed with the $\sigma$-algebra $\mathcal{C}$ generated by the corresponding evaluation maps $\{e_{t}:t\in I\}$. We also equip $\Gamma_{T}$ with the supremum distance and recall that the Borel $\sigma$-algebra generated by the corresponding open sets coincides with $\mathcal{C}$. We formulate the problem mentioned above as a variational problem on the space of path measures on $\Gamma_{T}$:
@@
$\mathrm{Lift}(\mu_{t})\coloneqq\Big{\{}\pi\in P(\Gamma_{T}):\quad(e_{t})_{\#}\pi=\mu_{t}\text{ for all }t\in I\Big{\}},$ (1.2)
where $e_{t}:\Gamma_{T}\to\mathcal{X}$ is the evaluation map defined by $e_{t}(\gamma)\coloneqq\gamma_{t}$ for $\gamma\in\Gamma_{T}$.
This extends Kantorovich’s classical problem, with two fixed marginals $\mu,\nu\in P(\mathcal{X})$, to a time-dependent setting with infinitely many fixed marginals $(\mu_{t})_{t\in I}\subset P(\mathcal{X})$. In contrast to the set of couplings $\mathrm{Cpl}(\mu,\nu)\subset P(\mathcal{X}^{2})$, which is always non-empty, the set of liftings $\mathrm{Lift}(\mu_{t})$ $\subset P(\Gamma_{T})\subset P(\mathcal{X}^{I})$ can be empty. In this case, the infimum (1.1) is +$\infty$+ ∞ by the usual convention.
@@
The reason for adopting a larger path space in the last case, namely the space of càdlàg curves $D(I;\mathcal{X})$, is that even $1$-absolutely continuous curves in 1-Wasserstein spaces, unlike the case $p>1$, cannot generally be lifted to measures on continuous paths. This can also be noticed through the Kolmogorov–Čentsov continuity criterion, where $p=1$ is excluded. In this work, we consider $p>1$ and study continuous Wasserstein curves with lower regularity than above.
@@
As the next step, we start from a path measure $\pi$ of finite $W^{\alpha,p}$-energy and show that its curve of one-dimensional time marginals $t\mapsto\mu_{t}\coloneqq(e_{t})_{\#}\pi$ inherits the same kind of regularity, and, moreover, its regularity is bounded from above by the energy of $\pi$. In fact, this transfer of regularity from the path measure to its curve of marginals holds for all lower-semi continuous norms listed in Table 1, as shown in Section 3.2, but is not reported here. Below is Theorem 3.5.
@@
By combining Proposition 1.1, Theorem 1.2, and the observation on $|\cdot|_{W^{\alpha,p}}$ in Table 1, we immediately draw the following conclusion as the first main result:
@@
    Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ can form a compatible collection under suitable conditions on their covariance matrices. See [19, page 49].
@@
    For more nontrivial compatible examples, see [7, Proposition 4.1] and [19, Section 2.3.2].
@@
    Construct the path measure $\pi_{n}\coloneqq(\ell)_{\#}\Upsilon_{n}\in P(\Gamma_{1})$, where $\ell:\mathcal{X}^{2^{n}+1}\to\Gamma_{1}$ is a $(\mathcal{B}(\mathcal{X}^{2^{n}+1})_{\Upsilon_{n}},\mathcal{C})$-measurable geodesic selection and interpolation map connecting the points with constant-speed geodesics. $\mathcal{B}(\mathcal{X}^{2^{n}+1})_{\Upsilon_{n}}$ denotes $\Upsilon_{n}$-completion of $\mathcal{B}(\mathcal{X}^{2^{n}+1})$.
@@
Notice that $\left(1-2^{-(p-\alpha p)}\right)\in(0,1)$ on the parameter range above. This characterization allows us to give the result below, which is a generalization of the metric Benamou–Brenier formula [5], [4, Eq. (8.0.3)], [14, Corollary 4], to the fractional setting. Below is Corollary 3.13 and $Geo([0,1];\mathcal{X})$ denotes the set of constant-speed geodesics.
@@ 1.2. Motivating example
In the first case of this example, we revisit the well-known relationship between the continuity equation and absolutely continuous Wasserstein curves under the $p$-integrability condition, [4, Chapter 8] and [14, Section 1]. In the second case, we assume a weaker integrability condition and highlight the estimates applied in the fractional setting.
@@
This can be used to infer that $(\mu_{t})$ is an absolutely continuous curve. Moreover, as shown in [14, Theorem 4] by the Lebesgue differentiation theorem, we obtain an upper bound for the metric speed $|\dot{\mu}_{t}|\leq\|\dot{\gamma}_{t}\|_{L^{p}(\pi)}$ for a.e. $t\in I$. Raising it to power $p$ and integrating in time yields
@@ 1.4. Acknowledgements
The author would like to thank his supervisor, Matthias Erbar, for invaluable guidance and numerous discussions. He also thanks Timo Schultz for insightful conversations, which significantly improved the results and eventually led to Examples 4.3, 4.4 and 4.5. He also thanks Zhenhao Li for helpful comments on some technical details. Additionally, he extends his thanks to Kohei Suzuki and Vitalii Konarovskyi for insightful discussions. The author acknowledges the financial support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.
@@ 2.1. Continuous paths on metric spaces
The Arzelà-Ascoli theorem provides a characterization of relative compactness in $\Gamma_{T}$ (see, e.g., [6, Theorem 7.2] or [10, Theorem 1.4]).
@@ 2.2. Hölder paths on metric spaces
Below, a simple yet useful characterization of Hölder continuity is provided. It states that for a continuous curve $X:[0,1]\to\mathcal{X}$ to be Hölder, it is enough to verify the Hölder condition only at the dyadic time points of $[0,1]$. The result is taken from [18, Lemma 2], whose proof is inspired by the classical proof of Kolmogorov–Čentsov continuity theorem (see e.g. [13, Theorem 2.8]).
Theorem 2.7 (A discrete characterization of $C^{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$ [18, Lemma 2]).
@@
While one direction is trivial, we stated the theorem in this way for its resemblance to the forthcoming Theorem 2.17. It is important to note that in the second condition, we evaluate $X$ only at a countable dense subset of the interval $[0,1]$. Thus, continuity needs to be added as an extra assumption. The result above differs slightly from the original statement in [18], where a (unique) continuous modification is constructed (for which one needs the space to be complete). Here due to the continuity assumption, the continuous modification coincides with the original curve at all times (and no additional assumption on the metric space is needed).
We recall some properties of the Hölder regularity. For below, see e.g. [10, Lemma 5.12].
@@ 2.3. $p$-variation paths on metric spaces
Two subsequent lemmas are taken from [10, Proposition 5.6].
@@
To distinguish these norms, we refer to the first one as variation, while the second one as infinitesimal variation, emphasizing that they are not necessarily identical. This is only the case for $p=1$ (see e.g. [10, Proposition 1.14]):
@@
As a next result, we recall the lower semi-continuity of $p$-variation, as stated in [10, Lemma 5.12].
@@
Let $(\mathcal{X},d)$ be a geodesic space. Take $p>1$ and let $\gamma$ be a continuous path such that $0<|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}<\infty$. Let $(D_{n})_{n\in\mathbb{N}}$ be a sequence of dissections of $[0,T]$ with shrinking mesh size $|D_{n}|\to 0$. Define $\gamma^{D_{n}}$ to be the piecewise geodesic approximation of $\gamma$ (i.e. the curve that coincides with $\gamma$ at points of $D_{n}$ and connects in between with geodesics). We know that $\gamma^{D_{n}}\to\gamma$ uniformly (see e.g. [10, Lemma 5.19]). For each $n$, the path $\gamma^{D_{n}}$ is of bounded $1$-variation, which implies, by Lemma 2.12, that $|\gamma^{D_{n}}|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}=0$ for all $p>1$. Consequently,
@@ 2.4. Fractional Sobolev paths on metric spaces
The fractional Sobolev spaces $W^{\alpha,p}$ (also known as Sobolev–Slobodeckij spaces) can be viewed as a specific instance of the broader class of Besov spaces $B^{\alpha,p,q}$ (see e.g. [16, Equation (2.1)]). In particular, when $q=p$, we have $B^{\alpha,p,p}=W^{\alpha,p}$. While the classical definition of $B^{\alpha,p,q}$ typically involves double integrals similar to the one mentioned above, there are also other function spaces whose definition involves double sums. Recently, [16] has proved that these definitions are in fact equivalent. For the purpose of our work, we only give their result for the case $q=p$:
Theorem 2.17 (A discrete characterization of $W^{\alpha,p}$ [16, Theorem 2.2]).
@@
This result is reminiscent of the discrete characterization of Hölder curves provided by [18, Lemma 2], as stated in Theorem 2.7. In the second definition, we are once again evaluating $X$ only at dyadic points (i.e. a countable subset of $[0,1]$) but this time, in the form of a double sum. Thus, continuity must be an additional assumption. This is, however, not needed for the first definition. In fact, under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the finiteness of $|X|_{W^{\alpha,p}}$ automatically implies the continuity of $X$ (see Theorem 2.19). We repeatedly rely on the equivalence result above. We proceed with the following simple observation:
@@
The estimate (2.29) below can be derived by applying the Garsia–Rodemich–Rumsey inequality, which was originally introduced in [11] for the case $\mathcal{X}=\mathbb{R}$. A proof for $\mathbb{R}^{\mathrm{d}}$ can be found in [22, Theorem 2.1.3] and for the general metric setting in [10, Theorem A.1]). The results below are well-known [10, Corollary A.2-3].
@@
Note that from (2.29), we can immediately conclude that $X$ has finite $({\alpha-1/p})^{-1}$-variation. But, as shown in [9, Theorem 2], by applying Hölder’s inequality with a clever choice of exponents, an even stronger statement can be made, namely, $X$ has finite $({\alpha})^{-1}$-variation. A consequence of this embedding is that under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the elements in $W^{\alpha,p}$ have zero infinitesimal $p$-variation by Lemma 2.12:
@@ 2.5. Sobolev paths on metric spaces
On $\mathbb{R}^{\mathrm{\,d}}$, it is well-known that the function spaces mentioned above coincide when $p>1$ (see e.g. [10, Proposition 1.45]). This result also holds in the general metric setting. For a proof in the case $p=2$, see [2, Theorem 10.2], which can be generalized to any $p>1$ and is stated below:
@@
which implies the embeddings [10, Theorem 1.47]
@@ 2.6. Geodesics on metric spaces
$d(X_{s},X_{t})=|t-s|d(X_{0},X_{1}),\quad\forall t,s\in[0,1].$